OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. c times the square root of c + 12 times c times the square root of c − 2 times c times the square root of c @jim_thompson5910

jimthompson5910 (jim_thompson5910):

let me try to write it out

OpenStudy (anonymous):

\[c \times \sqrt{c} + 12c \sqrt{c} - 2c \sqrt{c}\]

OpenStudy (anonymous):

There^

jimthompson5910 (jim_thompson5910):

oh you did it for me, thanks, ok one sec

jimthompson5910 (jim_thompson5910):

what you can do is replace every copy of \(\large \sqrt{c}\) with another variable, say z

OpenStudy (anonymous):

Okay.

jimthompson5910 (jim_thompson5910):

so \[\large c \times \sqrt{c} + 12c \sqrt{c} - 2c \sqrt{c}\] turns into \[\large cz + 12cz - 2cz\]

jimthompson5910 (jim_thompson5910):

what can you do from there?

OpenStudy (anonymous):

Ummmm. Solve? O.o

jimthompson5910 (jim_thompson5910):

see any like terms?

OpenStudy (anonymous):

Subtract 2cz from 12cz?

jimthompson5910 (jim_thompson5910):

good, what else

jimthompson5910 (jim_thompson5910):

cz + 12cz - 2cz is the same as 1cz + 12cz - 2cz

jimthompson5910 (jim_thompson5910):

if you subtract 2cz from 12cz, you get 10cz, so you now have 1cz + 10cz

OpenStudy (anonymous):

11cz

jimthompson5910 (jim_thompson5910):

we let \(\large z =\sqrt{c}\)

jimthompson5910 (jim_thompson5910):

so replace z with the square root of c to get this final answer \[\large 11c\sqrt{c}\]

OpenStudy (anonymous):

Awesome! Thanks so much.

jimthompson5910 (jim_thompson5910):

you're welcome